How do I solve by substitution?
2x+y=5
4x-3y=-5

Answer by Alan3354(69443)

2x+y=5 x3 --> 6x+3y= 15
4x-3y=-5
6x+3y= 15 Add eqns
---------
10x = 10
x = 1
------
Sub for x into either eqn
2*1 + y = 5
y = 3
------